OpenStudy (anonymous):

what is the solution of equations? 6x-2y=4 -9x+3y=12

OpenStudy (texaschic101):

are you familiar with the elimination method ?

OpenStudy (anonymous):

no @texaschic101

OpenStudy (anonymous):

please explain

OpenStudy (texaschic101):

6x - 2y = 4 9x + 3y = 12 what you are trying to do is eliminate a variable in the equation by making the coefficients cancel each other out. We can do this by multiplying. 6x - 2y = 4 -->(3) 9x + 3y = 12 -->(2) --------------- 18x - 6y = 12 (result of multiplying by 3) 18x + 6y = 24 (result of multiplying by 2) ----------------add 36x = 36 x = 1 you understand so far ?

OpenStudy (anonymous):

yess

OpenStudy (texaschic101):

now we are going to sub 1 in for x in either of the original equations 6x - 2y = 4 6(1) - 2y = 4 6 - 2y = 4 -2y = 4 - 6 -2y = -2 y = 1 now we can check our answers by subbing in known values into either of the original equations. 9x + 3y = 12 9(1) + 3(1) = 12 9 + 3 = 12 12 = 12 (correct) If the answer comes out equal, then it is correct. x = 1 and y = 1 or (1,1)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

no thanks

OpenStudy (texaschic101):

:)